OpenStudy (anonymous):

Evaluate and write the solution in scientific notation: 9.2 x 10^4 divided by 2.3 x 10^6. A. 4 x 10^2 B. 4 x 10^-2 C. 4 x 10^10

OpenStudy (anonymous):

\[9.2 \times 10^{4} \div 2.3 \times 10^{6} = 9.2 \div 2.3 \times 10^{4} \div 10^{6} = 4 \times 10^{-2}\]